OpenStudy (anonymous):

let y=(1/2)^x find the y intercept write an ordered pair and find the range either set or interval notation

OpenStudy (ranga):

To find the y-intercept, set x = 0 and solve for y.

OpenStudy (anonymous):

wouldn't y be 0 then also

OpenStudy (ranga):

No. (1/2)^0 = 1

OpenStudy (anonymous):

oh because the x is ^

OpenStudy (ranga):

correct.

OpenStudy (anonymous):

so (0,1) would be the ordered pair

OpenStudy (anonymous):

and that is the y intercept?

OpenStudy (ranga):

yes. That would be the ordered pair for the y-intercept.

OpenStudy (anonymous):

ok how do I find the Range then?

OpenStudy (ranga):

There is no restriction on x and therefore the domain for x is: (-infinity, infinity). If I put x = -very large number then y = (1/2)^(-very large number) 1 / 2^(-very large number) = 2^(very large number) When I bring the 2 to the numerator, the negative in the exponent goes away because 1/a^(-m) = a^m 2^(very large number) is a very large number and when the very large number is infinity then y will go to infinity too. positive infinity. When x becomes very large, then y becomes (1/2)^very large number. So the denominator will become very large and y will go to 0. So the range is (0, infinity)

OpenStudy (anonymous):

and on a graph I would start with the (0,1) and draw through the 0?

OpenStudy (anonymous):

or start at the (0,1) and draw up and towards the right on the graph?

OpenStudy (ranga):

It will look like this:

OpenStudy (anonymous):

ok thank you!

OpenStudy (ranga):

you are welcome.